The "Greasy Wrench" shop specializing in auto repairs rents a building at the. In the short run, it can vary its output (the number of cars fixed) only by varying the number of mechanics it employs. Each Greasy Wrench employee is paid a fixed wage.

Over the years, the owner of "Greasy Wrench", has tried hiring different numbers of workers. The data about the total output(Total product)for each number of workers are given in the table below

number of workers

1 Total product(TP)

2 5

3 14


4 30

44

50 Average Product(AP)

Marginal Product(MP)

54


5

6

Complete the table by calculating the Average Product(AP) of Labor and Marginal Product(MP) of Labor corresponding to each number of workers.

b. At which worker does the Law of Diminishing Marginal Returns start working? How do you know?

c. When Mr. Wei's son looked at the table, he said to his dad: "Hiring 5th and 6th person makes no sense!" Do you agree? Explain.


1
Expert's answer
2021-06-15T11:56:06-0400

a) Calculations:

APL = "\frac{TPL}{Laborers}", and MPL = TP of the next Worker - TP of the Preceding Worker .


So, APL of the;

1st worker = "\frac{5}{1}" = 5, MPL = 5 - 0 = 5


2nd worker ="\frac{14}{4}"= 7, and MPL = 14 - 5 = 9


3rd worker ="\frac{30}{3}"= 10, and MPL = 30 - 14 = 16


4th worker ="\frac{44}{4}"= 11, and MPL = 44-30 = 14


5th worker ="\frac{50}{5}"= 10, and MPL = 50 - 44 = 6


6th worker ="\frac{54}{6}"= 9, and MPL = 54 - 50 = 4


The above workings can be summarized in the table below:

No. of Workers Total Product Average Product(AP) Marginal Product

1 5 5 5

2 14 7 9

3 30 10 16

4 44 11 14

5 50 10 6

6 54 9 4


b) The Law of Diminishing Marginal Returns start working at the 4th, the Marginal Product of Labor starts to reduce (diminish) from 16 for the 3rd worker to 14 (for the 4th worker).


c) I agree with Mr. Wei's view; the productivity of labor is optimized at the 3rd worker (MPL is 16) while it diminishes at the 4th worker (MPL of 14) though it’s still economical, the employment of the 5th (MPL = 6), and 6th (MPL = 4) worker will constrain the shop's revenue while bringing little additional benefit. The 6th worker would be much uneconomical for the enterprise.